A pendulum is formed from a small rock with mass of 0.10 kg that is fastened to a massless string




with length 0.750 m. It is swinging so as to make a maximum angle of 45.0 with the vertical. Air




resistance is negligible. What is the tension in the string as it passes through the vertical?

1
Expert's answer
2022-04-17T17:06:58-0400

The tension in the string as pendulum passes through the vertical

"T=mg+\frac{mv^2}{l}"

The speed of rock at this instant

"v=\sqrt{2gl(1-cos\theta)}"

Hence

"T=mg+2mg(1-cos\theta)=mg(3-\cos\theta)\\
=0.10*9.8*(3-\cos 45^\circ)=2.25\:\rm N"